If a(b-c)x^2 + b(c-a) x + c(a-b) =0 hai equal roots then 2/b=

Grade:12

2 Answers

Saurabh Koranglekar
5 years ago
Dear student

D = 0

b^2 -4ac = 0

using the quadratic formula we can find the required
Aditya Gupta
5 years ago
saurabh’s answer is sonce again not feasible.
let f(x)= a(b-c)x^2 + b(c-a) x + c(a-b) 
note that f(1)=0
that means both roots should be 1.
so 1+1= -b(c-a)/a(b-c)
and 1*1= c(a-b)/a(b-c)
from these eqns it is much easier to find b